# Does Positive Reinforcement Work?: A Quasi-Experimental Study of the Effects of Positive Feedback on Reddit

### v1

@@ 2.4. Gift-Giving and Reciprocity
Social computing literature highlights gift-giving as a mechanism that builds a sense of reciprocity from gift-givers. Essentially, users expect to receive similar treatment from their communities as they give (Kollock et al., 1999). Other work shows that receiving a gift encourages users to give similar gifts in the future (Kizilcec et al., 2018), highlighting reciprocity for the recipient. Importantly, gift-giving is a means to build reputation which social computing literature touts as, a clear signal of desired behaviors (Lampe, 2012), an incentive for good behavior (Resnick et al., 2000), and a tool to build trust and engagement among users (Resnick et al., 2006).
In Reddit communities, positive feedback mechanisms are staples of gift-giving, and if they abide by the principles of reciprocity, recipients of positive feedback may be motivated to provide similar feedback in the future. This highlights the importance of understanding user-level outcomes of gift-giving through positive feedback. Our work explores the user-level outcomes in response gift-giving in the form of Reddit gilds and upvotes.
@@ 3. Study Design and Data
Figure 2. Pipeline to go from data collection to groups of matched treated and control samples used in our causal inference methodology.
@@
Pipeline diagram of our data collection and analysis, split into three boxes. The first box indicates the collection of the dataset from Pushshift’s archives of the top 100 subreddits that gild most frequently, resulting in 10.9M posts. The second box shows how we identify a pool of candidates from this large initial sample. After various filters we have 13.4K treated and 40.3K control candidates for the gild treatment. For the score treatment, we have 266K treated and 548K control candidates. The final box depicts our process for performing stratified matching and eventually results in 85 strata made of 4.6K treated and 14.7K control users for the gild treatment. The score treatment is left with 646 strata consisting of 258K treated and 480K control users.
This section describes our data-collection process and the causal-inference approach we adopted drawing on the potential outcome framework (Rosenbaum and Rubin, 1983; Rubin, 2005) used in other social computing research (Jhaver et al., 2024; Chandrasekharan et al., 2017a; Saha et al., 2019a). Our approach involves identifying a treatment group (posts that receive positive feedback) and a group of control users (posts that did not receive positive feedback). We identify similar posts from each group to create matches, allowing us to compare the trajectory of similar users who differ only in whether they received our treatment (i.e., positive feedback). The full pipeline of this process is visualized in Figure 2 along with descriptive statistics about the dataset.
@@ 3.1. Treatments: Posts That Received Positive Feedback
There are many forms of feedback available on Reddit. We select two to focus on in this work: gilds and score. Figure 1 visualizes both forms of feedback in a real Reddit post.
Gilds were an award-like signal consisting of gold icons that were awarded to posts. Gilds and other awards were discontinued by Reddit in 2023, [^6] https://www.reddit.com/r/reddit/comments/14ytp7s/reworking_awarding_changes_to_awards_coins_and/ however gilds were considered beloved by some users on Reddit (Weatherbed, 2024) and many users were disappointed that they were not among the features brought back to the platform in May 2024. Thus, we believe there is still value in exploring how gilds affected the recipients to understand whether Reddit removed a mechanism that had positive impacts on user behavior and experience.
Because gilds are a paid feature, they are a potentially strong signal of quality given the rarity with which they were given out. However, the real-world cost of gilding means they may not be entirely representative of all opinions or values from a community. To account for this, we select score (i.e., aggregated upvotes and downvotes) as our second treatment. Upvotes are the most widespread signal of approval on Reddit, utilized by every community. To contrast gilds, upvotes are free to give out, therefore there is no limit on how many a user can hand out, neither explicitly imposed by the platform nor implicitly imposed by cost. We note that some moderators believe that upvotes are not necessarily a reliable indicator of quality, specifically in the context of communities like r/AskHistorians that value input from experts (Jiang et al., 2023). However, these moderators are not claiming that upvotes cannot reliably inform us what a community likes. Thus, we recognize this limitation of our work and do not claim that the posts treated with upvotes contain accurate information or even content that moderators themselves would like to encourage. We rely on the concept of contingency (Miltenberger, 2016) by assuming that receiving the treatments is contingent upon posting something deserving of reinforcement within the community. Thus, we trust community signals of approval and do not evaluate how deserving the treated posts are. This mimics approaches of prior work which focus only on the outcomes of treatment and do not consider the precursors leading to the treatment being applied (Jhaver et al., 2019b, 2024).
@@ 3.2. Identify Pool of Candidates
Since upvotes are used universally and gilds are not utilized by all communities, we construct a sample from the 100 subreddits which awarded the most gilds in the study period. We form distinct pools of treatment and control posts, $\mathcal{C}_{\text{gild}}$ and $\mathcal{C}_{\text{score}}$, for each treatment, made within our treatment window: May 15, 2020 to September 17, 2020. This smaller window allows us to study user behavior in the 14 days before and after treatment. Posts which received at least one gild are considered treated candidates in $\mathcal{C}_{\text{gild}}$. Posts that receive no gilds are added to $\mathcal{C}_{\text{gild}}$ as control candidates.
In $\mathcal{C}_{\text{score}}$, however, we need to identify some threshold above which a post is considered to be highly upvoted. To accomplish this, we split the distribution of scores for all posts in each subreddit into 4 quantiles per month. The highest quantile consists of the highest-scored posts which are added to $\mathcal{C}_{\text{score}}$ as treated posts. Posts in the bottom two quantiles comprise the control candidates in $\mathcal{C}_{\text{score}}$. We drop all posts in the second highest quantile to provide a buffer between control scores and treated scores and to ensure that our treatment is strongly defined.
For each treated and control candidate in $\mathcal{C}_{\text{gild}}$ and $\mathcal{C}_{\text{score}}$, we ensure that the author of the post has made at least one post in the 14 days prior, none of which received the respective treatment. This provides us with baseline activity levels and avoids users receiving multiple treatments in a short span of time. In this setup, it is possible for one user to appear in both candidate pools multiple times, a fluid interpretation of our treatments since the start of our study period is arbitrary so we do not know which users receive treatment before the start of our analysis.
For both treatments, we sample 3 times as many control users from each subreddit as there are treated users to balance the groups while still providing sufficient control candidates for matching. Figure 2 reports summary statistics about the final pools of candidates, $\mathcal{C}_{\text{gild}}$ and $\mathcal{C}_{\text{score}}$.
For each candidate in $\mathcal{C}_{\text{gild}}$ and $\mathcal{C}_{\text{score}}$, we dynamically define three events of interest:
@@
  • Treatment/placebo time: the creation time of the post that was treated or received the placebo.
@@
We note that our analysis is limited by not knowing the precise time when treatment was applied, nor do we have confirmation that the author of the treated post saw that they received the positive feedback. As a result, we use the creation time of the treated post as a proxy for treatment time. This may raise issues with immediacy (Miltenberger, 2016) in that authors are likely not treated immediately after contributing and they may not realize they received the treatment until even later. As a result, our analysis is a conservative estimate of the effect of positive feedback on users.
@@ 3.3. Stratified Matching
We want to minimize the differences in pre-treatment author and environmental characteristics between the control and treatment groups so observed effects can be attributed to the treatments. For each treated user, our goal to identify similar control users from the same community based on pre-treatment posting behavior and content quality.
We perform within-subreddit stratified matching on propensity scores within each candidate pool. Stratified propensity score matching has been used in prior work (Kiciman et al., 2018; Saha and Sharma, 2020; Saha et al., 2018; Verma et al., 2022; Yuan et al., 2023; Olteanu et al., 2017) to minimize biases of propensity score matching identified by King and Nielsen (2019). For each subreddit, we train an AdaBoost classifier which outputs a propensity score given the following covariates (log-scaled to achieve more normal distributions) as independent variables:
@@
  • Frequency of the 100 most common bigrams from the titles and post bodies of all posts in our dataset.
@@
  • Average number of gilds received per post in the baseline period.
This variable does not apply to authors of posts in $\mathcal{C}_{\text{gild}}$ because, by definition, they must not have been gilded in their baseline period.
The resulting propensity scores reflect how likely a user is to belong to the treatment group based on the covariates. Within each subreddit, we bin the propensity scores into 10 strata using quantiles, filter out the strata with fewer than 10 users, and ensure that each stratum contains both treated and control users.
@@
A plot with two sub-plots. The leftmost plot reports the average Standardized Mean Difference (SMD) of each covariate for the matched data using the gild treatment. The plot shows that the unmatched data has higher SMD values than the matched data for all covariates except the N-grams. A dotted line at the SMD value of 0.3 shows that all matched data has SMD values below our threshold of 0.3. The rightmost plot visualizes the same measures for the matched data using the score treatment. All covariates except the average number of gilds has a lower SMD after matching, however no matched data surpasses the line at 0.3 SMD.
To measure the quality of our matches, we compute the Standardized Mean Difference (SMD) between the treatment and control samples of each stratum for each covariate. We use an upper bound of 0.3 to indicate good match quality overall (Kiciman et al., 2018; Kampenes et al., 2007). We drop all strata with a mean SMD over all covariates above 0.3 to ensure that we retain only well-matched treatment and control posts to obtain an unbiased estimate of the treatment effects. Figure 3 visualizes the SMD values for each of our covariates averaged across the remaining strata (i.e., matched samples) and across the original pool of candidates (i.e., unmatched samples) to demonstrate that matching improved the SMDs for our covariates. We note that the SMDs for two of the covariates were slightly lower before matching, however the mean matched SMDs are still below our threshold of 0.3. We report descriptive statistics about the final matches in Figure 2.
3.4. Outcomes
With this dataset of well-matched treatment and control posts, we examine the effects of receiving positive feedback on three major user-level themes: reception from the community, engagement, and norm-adherence. To measure community reception, we look at the effect of our treatments on each user’s future average number of gilds and future average score per day. These outcomes capture whether receiving positive feedback encourages users to post more positive-feedback-worthy content in the future. For engagement, prior work describes positive feedback as a mechanism capable of increasing user motivation to participate Kraut et al. (2011). We use daily posting rate as a proxy for user motivation to participate. Finally, to capture the impact of treatment on norm-adherence, we utilize moderation decisions by each community’s centralized moderation team through each user’s daily removal rate.
For each post in our matched sample, we collect all other posts by the same author in the baseline and observation windows. This gives us a set of 309K posts for our gild treatment and 5.4M posts for our score treatment. We calculate daily values for each of our four outcomes, taking care to exclude the treated or placebo post to not skew any averages. This results in 129K daily aggregates for the gild treatment and 2.6M daily aggregates for the score treatment.
Figure 4. Visualization of each outcome over the baseline and observation windows (excluding the treatment/placebo posts). These plots are used to validate the parallel trends assumption needed for difference-in-differences analysis.
[…]
This figure has eight subplots. The top row of four plots visualize each outcome over time for the gilded posts and their matches. The bottom four subplots visualize the same outcomes for the highly-upvoted posts and their matches. The plots start at the beginning of the baseline period (day -14) and extend to the end of the observation window (day 13). These plots show that most outcomes are parallel prior to treatment between the treatment and control groups, but the treatment groups experience different effects after treatment compared to the control groups.
@@ 4. RQ1: What is the effect of positive feedback on recipients?
4.1. RQ1a: How does receiving a gild or being highly upvoted impact users’ future behavior in a community?
To understand how the treatments affect users (RQ1a), we conduct a Difference-in-Differences (DiD) analysis similar to analyses done in prior work (Chandrasekharan et al., 2017a; Abadie, 2005; Grinberg et al., 2016) which allows us to control for any time-related effects on our outcomes. For example, users in smaller communities experiencing a surge in subscribers may receive higher score on their posts across the board. The DiD analysis compares the difference in treated user’s behavior before and after receiving treatment against the difference the control group experiences over time.
To estimate causal effects through DiD analysis, the data must satisfy the parallel trends assumption (did, 2016). This means the difference between the outcomes for the treatment and control groups must be constant over time, prior to treatment. Similar to prior work (Marcus and Sant’Anna, 2021; Ryan et al., 2019), we visually inspect our data in Figure 4 and conclude that the parallel trends assumption is satisfied in all but two cases, daily post rate and daily removal rate for the gild outcome, which we exclude from the DiD analysis. This ensures the internal validity of the DiD models for the other outcomes and any observed effects (did, 2016).

### v2

@@ 2.4. Gift-Giving and Reciprocity
Social computing literature highlights gift-giving as a mechanism that builds a sense of reciprocity from gift-givers. Essentially, users expect to receive similar treatment from their communities as they give (Kollock et al., 1999). Other work shows that receiving a gift encourages users to give similar gifts in the future (Kizilcec et al., 2018), highlighting reciprocity for the recipient. Importantly, gift-giving is a means to build reputation which social computing literature touts as a clear signal of desired behaviors (Lampe, 2012), an incentive for good behavior (Resnick et al., 2000), and a tool to build trust and engagement among users (Resnick et al., 2006).
In Reddit communities, positive feedback mechanisms are staples of gift-giving, and if they abide by the principles of reciprocity, recipients of positive feedback may be motivated to provide similar feedback in the future. This highlights the importance of understanding user-level outcomes of gift-giving through positive feedback. Our work explores the user-level outcomes in response gift-giving in the form of Reddit gold and upvotes.
@@ 3. Study Design and Data
Figure 2. Pipeline from data collection to forming matched treated and control samples used in our causal inference methodology.
@@
Pipeline diagram of our data collection and analysis, split into three boxes. The first box indicates the collection of the dataset from Pushshift’s archives of the top 100 subreddits that gild most frequently, resulting in 10.9M posts. The second box shows how we identify a pool of candidates from this large initial sample. After various filters we have 13.4K treated and 40.3K control candidates for the gold treatment. For the score treatment, we have 266K treated and 548K control candidates. The final box depicts our process for performing stratified matching and eventually results in 85 strata made of 4.6K treated and 14.7K control users for the gold treatment. The score treatment is left with 646 strata consisting of 258K treated and 480K control users.
This section describes our data-collection process and the causal-inference approach we adopted drawing on the potential outcomes framework (Rosenbaum and Rubin, 1983; Rubin, 2005) used in prior social computing research (Jhaver et al., 2024; Chandrasekharan et al., 2017a; Saha et al., 2019a). Our goal is to examine the causal effects of receiving positive feedback (i.e., the treatment) compared to what would have happened if the treatment was not administered (i.e., the counterfactual). However, it is not possible to find a post that simultaneously did and did not receive positive feedback, so a true counterfactual is missing or unobserved. The potential outcomes framework allows us to estimate the missing counterfactual for each treated post based on the outcomes of similar, observed posts that were not treated (i.e., did not receive positive feedback). We estimate the counterfactual by identifying a treatment group (posts that receive positive feedback) and a similar control group (posts that did not receive positive feedback). We identify similar posts from each group to create matches, allowing us to compare the trajectory of similar users who differ only in whether they received our treatment (i.e., positive feedback). The full pipeline of this process is visualized in Figure 2 along with descriptive statistics about the dataset.
@@ 3.1. Treatments: Posts That Received Positive Feedback
There are many forms of feedback available on Reddit. We select two to focus on in this work: gold and score. Figure 1 visualizes both forms of feedback in a real Reddit post.
Gold was an award-like signal consisting of gold icons that were awarded to posts. Gold and other awards were discontinued by Reddit in 2023, [^7] https://www.reddit.com/r/reddit/comments/14ytp7s/reworking_awarding_changes_to_awards_coins_and/ however gold was considered beloved by some users on Reddit (Weatherbed, 2024) and many users were disappointed that they were not among the features brought back to the platform in May 2024. Thus, we believe there is still value in exploring how gold affected the recipients to understand whether Reddit removed a mechanism that had positive impacts on user behavior and experience.
Because gold was a paid feature, it is a potentially strong signal of quality given the rarity with which it was given out. However, the real-world cost of gilding means that gold may not be entirely representative of all opinions or values from a community. To account for this, we select score (i.e., aggregated upvotes and downvotes) as our second treatment. Upvotes are the most widespread signal of approval on Reddit, utilized by every community. To contrast gold, upvotes are free to give out, therefore there is no limit on how many a user can hand out, neither explicitly imposed by the platform nor implicitly imposed by cost. We note that some moderators believe that upvotes are not necessarily a reliable indicator of quality, specifically in the context of communities like r/AskHistorians that value input from experts (Jiang et al., 2023). However, these moderators are not claiming that upvotes cannot reliably inform us what a community likes. Thus, we recognize this limitation of our work and do not claim that the posts treated with upvotes contain accurate information or even content that moderators themselves would like to encourage. Furthermore, we rely on the concept of contingency (Miltenberger, 2016) by assuming that receiving the treatments is contingent upon posting something deserving of reinforcement within the community. Thus, we trust community signals of approval and do not evaluate how deserving the treated posts are. This mimics approaches of prior work which focus only on the outcomes of treatment and do not consider the precursors leading to the treatment being applied (Jhaver et al., 2019b, 2024).
@@ 3.2. Identify Pool of Candidates
Since upvotes are used universally and gold was not utilized by all communities, we construct a sample from the 100 subreddits which awarded the most gold in the study period. We form distinct pools of treatment and control posts, $\mathcal{C}_{\text{gold}}$ and $\mathcal{C}_{\text{score}}$, for each treatment, made within our treatment window: May 15, 2020 to September 17, 2020. This smaller window allows us to study user behavior in the 14 days before and after treatment. Posts which received at least one gold are considered treated candidates in $\mathcal{C}_{\text{gold}}$. Posts that receive no gold are added to $\mathcal{C}_{\text{gold}}$ as control candidates.
To construct $\mathcal{C}_{\text{score}}$, we need to identify some threshold above which a post is considered to be highly upvoted. To accomplish this, we split the distribution of scores for all posts in each subreddit into 4 quantiles per month. The highest quantile consists of the highest-scored posts which are added to $\mathcal{C}_{\text{score}}$ as treated candidates. Posts in the bottom two quantiles comprise the control candidates in $\mathcal{C}_{\text{score}}$. We drop all posts in the second highest quantile to provide a buffer between control scores and treated scores, ensuring that our treatment is strongly defined.
[…]
A plot with two sub-plots. The leftmost plot reports the average Standardized Mean Difference (SMD) of each covariate for the matched data using the gold treatment. The plot shows that the unmatched data has higher SMD values than the matched data for all covariates except the N-grams. A dotted line at the SMD value of 0.3 shows that all matched data has SMD values below our threshold of 0.3. The rightmost plot visualizes the same measures for the matched data using the score treatment. All covariates except the average number of gold has a lower SMD after matching, however no matched data surpasses the line at 0.3 SMD.
For each treated and control candidate in $\mathcal{C}_{\text{gold}}$ and $\mathcal{C}_{\text{score}}$, we ensure that the author of the post has made at least one post in the 14 days prior, none of which received the respective treatment. This provides us with baseline activity levels and avoids users receiving multiple treatments in a short span of time. In this setup, it is possible for one user to appear in both candidate pools multiple times, a fluid interpretation of our treatments since the start of our study period is arbitrary so we do not know which users receive treatment before the start of our analysis.
For both treatments, we sample 3 times as many control users from each subreddit as there are treated users to balance the groups while still providing sufficient control candidates for matching. Figure 2 reports summary statistics about the final pools of candidates, $\mathcal{C}_{\text{gold}}$ and $\mathcal{C}_{\text{score}}$.
For each candidate in $\mathcal{C}_{\text{gold}}$ and $\mathcal{C}_{\text{score}}$, we dynamically define three events of interest:
@@
  • Treatment/placebo time: the creation time of the post that received the treatment or placebo.
@@
We note that our analysis is limited by the fact that the precise time at which positive feedback is given is not publicly accessible, so we do not have confirmation that the recipients are aware that the treatment has been applied. We use the creation time of the treated post as a proxy for treatment time which may raise issues with immediacy (Miltenberger, 2016), in that authors are likely not treated immediately after contributing and they may not realize they received the treatment until even later. We could have further restricted the treatment pool to authors who also made comments on their gilded post, however it is not guaranteed that the gold was given prior to their comment. Thus, we consider our analysis to be a conservative estimate of the effect of positive feedback on users.
@@ 3.3. Stratified Matching
We want to minimize the differences in pre-treatment author and environmental characteristics between the control and treatment groups so observed effects can be attributed to the treatments. For each treated user, our goal is to identify similar control users from the same community based on treatment or placebo post content and pre-treatment posting behavior and content quality.
We perform within-subreddit stratified matching on propensity scores in each candidate pool. Stratified propensity score matching has been used in prior work (Kiciman et al., 2018; Saha and Sharma, 2020; Saha et al., 2018; Verma et al., 2022; Yuan et al., 2023; Olteanu et al., 2017) to minimize biases of propensity score matching identified by King and Nielsen (2019). This approach allows us to handle the bias-variance trade-off by striking a balance between too biased (one-to-one matched) and too variant (unmatched) data comparisons so we can isolate and examine the effects of treatment within each stratum (Jhaver et al., 2024; Saha et al., 2021). For each subreddit, we train an AdaBoost classifier which outputs a propensity score given the following covariates (log-scaled to achieve more normal distributions) as independent variables:
@@
  • Frequency of 100 bigrams from the text contained in the treatment/placebo post (i.e., titles and post bodies). We use the 100 most common bigrams appearing in all of the posts in $\mathcal{C}_{\text{gold}}$ or $\mathcal{C}_{\text{score}}$.
@@
  • Average number of gold received per post in the baseline period.
This variable does not apply to authors of posts in $\mathcal{C}_{\text{gold}}$ because, by definition, they must not have been gilded in their baseline period.
The resulting propensity scores reflect how likely a user is to belong to the treatment group based on the covariates. These covariates allow us to match treatment and control users that have similar “quality” posts in the pre-treatment window, so the authors should be similarly good writers before the treatment occurs. After matching, any effects observed after the treatment is applied can be attributed to the treatment itself. Within each subreddit, we bin the propensity scores into 10 strata using quantiles, filter out the strata with fewer than 10 users, and ensure that each stratum contains both treated and control users.
@@
To measure the quality of our matches, we compute the Standardized Mean Difference (SMD) between the treatment and control samples of each stratum for each covariate. We use an upper bound of 0.3 to indicate good match quality overall (Kiciman et al., 2018; Kampenes et al., 2007). We drop all strata with a mean SMD over all covariates that exceeds 0.3 to ensure that we retain only well-matched treatment and control posts and obtain an unbiased estimate of the treatment effects. Figure 3 visualizes the SMD values for each of our covariates averaged across the remaining strata (i.e., matched samples) and across the original pool of candidates (i.e., unmatched samples) to show that matching improved the SMDs for our covariates. We note that the SMDs for two of the covariates were slightly lower before matching, however the mean matched SMDs are still below our threshold of 0.3. We report descriptive statistics about the final matches in Figure 2.
3.4. Outcomes
With this dataset of well-matched treatment and control posts, we examine the effects of receiving positive feedback on four major user-level themes: reception from the community, content, engagement, and norm-adherence. To measure community reception, we look at the effect of our treatments on each user’s future average number of gold and future average score per day. These outcomes capture whether receiving positive feedback encourages users to post more positive-feedback-worthy content in the future. To understand the effect of positive feedback on content, we measure the text difference between the treatment or placebo post and the recipient’s future contributions. This difference is measured by computing sentence embeddings (Reimers and Gurevych, 2019) for the text in each post’s title and body and measuring the cosine distance to the text of their treated or placebo post. For engagement, prior work describes positive feedback as a mechanism capable of increasing user motivation to participate Kraut et al. (2011). We use daily posting rate as a proxy for user motivation to participate. Finally, to capture the impact of treatment on norm-adherence, we utilize moderation decisions by each community’s centralized moderation team through each user’s daily removal rate.
For each post in our matched sample, we collect all other posts by the same author in the baseline and observation windows. This gives us a set of 309K posts for our gold treatment and 5.4M posts for our score treatment. We calculate daily values for each of our five outcomes, taking care to exclude the treated or placebo post to not skew any averages. This results in 129K daily aggregates for the gold treatment and 2.6M daily aggregates for the score treatment.
Figure 4. Visualization of each outcome over the baseline and observation windows (excluding treatment/placebo posts themselves). These plots are used to validate the parallel trends assumption needed for difference-in-differences analysis.
@@ 3.4. Outcomes
This figure has ten subplots. The top row of five plots visualize each outcome over time for the gilded posts and their matches. The bottom five subplots visualize the same outcomes for the highly-upvoted posts and their matches. The plots start at the beginning of the baseline period (day -14) and extend to the end of the observation window (day 13). These plots show that most outcomes are parallel prior to treatment between the treatment and control groups, but the treatment groups experience different effects after treatment compared to the control groups.
@@ 4. RQ1: What is the effect of positive feedback on recipients?
Using our matched treatment and control samples, we perform analyses to understand the causal impact of receiving positive feedback on user behavior and how long those effects last.
4.1. RQ1a: How does receiving gold or being highly upvoted impact users’ future behavior in a community?
To understand how the treatments affect users (RQ1a), we conduct a Difference-in-Differences (DiD) analysis similar to analyses done in prior work (Chandrasekharan et al., 2017a; Abadie, 2005; Grinberg et al., 2016). The DiD analysis compares the difference in treated user’s behavior before and after receiving treatment against the difference the control group experiences over time. Like any causal inference study using observational data, there are limitations to our approach (Weld et al., 2022). Our matching process accounts for measured confounds in our data (i.e., the covariates) and the DiD analysis captures time-invariant unmeasured confounds, for example, users in small communities experiencing a surge in subscribers may receive higher score on their posts across the board. It is possible, however, that external events impact how a community reacts to content in ways not captured by our observational data, leaving some time-variant unmeasured confounds unaccounted for. Future work can employ true experiments to account for such confounds and extend our understanding of the impact of positive feedback.
To estimate causal effects through DiD analysis, the data must satisfy the parallel trends assumption (did, 2016). This means the difference between the outcomes for the treatment and control groups must be constant over time, prior to treatment. Similar to prior work (Marcus and Sant’Anna, 2021; Ryan et al., 2019), we visually inspect our data in Figure 4 and conclude that the parallel trends assumption is satisfied in all but two cases, daily post rate and daily removal rate for the gold outcome, which we exclude from the DiD analysis. This ensures the internal validity of the DiD models for the other outcomes and any observed effects (did, 2016).